OpenStudy (anonymous):

demonstrate how your quadratic equation expression is solved by the Quadratic Formula. Explain your steps x²+8x+2=0

OpenStudy (anonymous):

You need to factor

OpenStudy (anonymous):

Solutions are: -.25 and -7.74

OpenStudy (anonymous):

Using the quadratic formula which is \[(-b \pm {\sqrt{b^2-4ac}})/2a\]

OpenStudy (anonymous):

Actually mridrik is right, you should use the quadratic formula

OpenStudy (anonymous):

the opposite of b (8) = -8

OpenStudy (anonymous):

plus or minus sqrt(b^2-4ac) the square root of (8^2-4(2)(1)) = sqrt(56)

OpenStudy (anonymous):

then divide by 2

OpenStudy (anonymous):

your two solutions are 8+sqrt(56) = 8 + 2sqrt(14) divided by two is 4 + sqrt(14) since we have to do plus or minus the other solution is 4 - sqrt(14)